the ip of h is 13.6 ev. it is exposed to electromagnetic waves of 1028A and given out induced radiations. find the wavelength of these induced radiations

Grade:11

1 Answers

Arun
5 years ago
Dear Kabir
 

H-atom is exposed to UV radiation having wavelength 102.8nm

 

the energy of photon corresponding to 102.8 nm waveegth is

 

E = hc/λ = 1242 eV-nm /102.8 nm

 

   = 12.42 eV

 

The energy needed to take athe H atom frm ground state to first excited state is

 

                    E2-E1 - 13.6 eV - 3.4 eV = 10.2 eV

 

and to  the second excited state is

 

                    E3-E1 - 13.6 eV - 1.5 eV = 12.1 eV

 

and to  the third excited state is

 

                    E4-E1 - 13.6 eV - 0.85 eV = 12.75 eV

 

thus 12.42 eV photon has large probability of being absorbed in 12.1 eV

 

The corresponding  waveegth is: λ =103 nm

 

Regards

Arun